OpenStudy (anonymous):

The length of a rectangle is 3 inches less than twice the width. If the perimeter of the rectangle is 18 inches, find the width of the rectangle.

OpenStudy (yanasidlinskiy):

The length of a rectangle is 3 inches less than twice the width Let x and (2x-3) represent the width and length respectively 2(2x-3) + 2x = 18in 6x = 24in x = 4in, the width. The length is 5in (2*4in-3in) CHECKING our Answer*** 2*5in + 2*4in = 18in

OpenStudy (anonymous):

The length of a rectangle is 3 inches less than twice the width Let x and (2x-3) represent the width and length respectively 2(2x-3) + 2x = 18in 6x = 24in x = 4in, the width. The length is 5in (2*4in-3in) CHECKING our Answer*** 2*5in + 2*4in = 18in

OpenStudy (texaschic101):

de ja vous.....I am seeing double....lol